# MFCPL: Multimodal Federated Cross Prototype Learning for Missing Modalities and Data Heterogeneity

### v1

@@ 4.1 Experimental Setup
In the case of modalities such as image, text and audio, we follow Feng et al. (2023) to extract features for local training. This method involves deriving features instead of training from scratch using raw data, a choice made to alleviate the computational burden. Our model architecture is composed of three key modules: the encoder, the fusion layer, and the classifier. For audio, accelerometer, and gyroscope data, the encoder adopts a 1D Conv+RNN architecture. The 1D Conv includes three CNN layers followed by two MLP layers, while the RNN component incorporates a single GRU layer. In contrast, the encoder for images is built with two MLP layers, and for text, we employ an RNN architecture with a single GRU layer. The fusion layer is implemented as a self-attention fusion layer, and the classifier is subsequently followed by two MLP layers. The hidden size for GRU layer is set as $128$ for all datasets. Regarding the projection head, we utilize a single MLP layer with the default dimension set to $64$.
Implementation Details.
@@
Following the methodology outlined in Feng et al. (2023), we set a uniform missing rate of $q=\{0.5,0.7,0.8\}$ for each modality based on Bernoulli distribution, creating a scenario that simulates the presence of missing modalities. The communication round is set to $200$, and the local training epoch is $1$ for all datasets. The default value for the temperature $\tau$ is $0.1$. The batch size is set to $16$, and we employ the SGD Robbins and Monro (1951) optimizer with a weight decay of $51e-$ 5, maintaining a consistent learning rate of $0.05$ across all datasets. We tunedhyperparameters $\alpha_{reg}$, $\alpha_{con}$, and $\alpha_{align}$ within the range of $\{0.1,3\}$ and configured with values yielding the best results for each dataset. For the Hateful Memes dataset, we use the AUC score as the evaluation metric. The UAR score serves as the metric for the MELD dataset, and the F1 score is employed as the metric for the UCI-HAR dataset. We repeat the experiment $5$ times and report mean values.
Baselines.
@@
For evaluation, we compare our MFCPL with five baselines in MFL setting: (1) FedAvg McMahan et al. (2017), (2) FedProx Li et al. (2020), (3) FedOpt Reddi et al. (2020), (4) MOON Li et al. (2021), (5) FedProto Tan et al. (2022b) (with parameter averaging).
@@
Figure 3: Illustrative example of average F1 score versus communication rounds on UCI-HAR dataset with $q=0.5$.
@@
q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8
70.90 69.46 67.50 54.56 50.09 49.56 51.74 50.71 50.58
75.09 74.31 73.63 53.96 54.92 54.15 52.16 51.69 49.98
71.05 70.55 71.68 51.42 51.58 50.23 51.83 50.45 50.43
73.56 70.96 70.41 51.79 51.17 50.63 51.77 52.05 51.03
77.44 75.61 75.19 56.30 55.69 55.04 53.73 52.24 51.92
4.2 Experimental Results
Performance Comparison.
Table 1 illustrates the performance comparison of our proposed MFCPL with other state-of-the-art methods on three multimodal federated datasets. Our method, MFCPL, consistently outperforms other baselines across various missing rates, highlighting its effectiveness in achieving better generalized global models and supporting clients with missing data. As the missing rate parameter $q$ increases, MFCPL maintains its performance, while other baselines exhibit a significant decrease in accuracy. Taking the UCI-HAR dataset with $q=0.5$ as an illustrative example, our method surpasses the best-performing baseline FedProto by a notable margin, showcasing an improvement of $2.57\%$. These results underscore the robustness and effectiveness of MFCPL in handling missing data scenarios across different client types and datasets. Figure 3 depics the performance curves (F1 score versus communication rounds) of our MFCPL and SOTA baselines, from which we can observe that MFCPL demonstrates a more stable convergence compared to other methods, particularly under non-IID and missing modality conditions.
Prototype UCI-HAR Hateful Memes MELD
Unimodal Prototypes 77.09 54.95 52.13
Complete Prototypes 77.44 56.30 53.73
Table 3: Performance analysis of MFCPL on utilizing complete prototypes and unimodal prototypes with $q=0.5$.
@@ 4.3 Comparison to Unimodal Prototypes
To assess the efficacy of our proposed complete prototypes, we conducted experiments comparing them with unimodal prototypes using $q=0.5$, as summarized in Table 3. The results demonstrate the superior performance of utilizing our proposed complete prototypes compared to the unimodal prototypes. Notably, our method exhibits a performance improvement of $1.6\%$ on the MELD dataset over the unimodal prototypes. This finding highlights the effectiveness of capturing diverse modality knowledge from complete prototypes, providing valuable guidance for the training process.
@@ 4.4 Ablation Studies
Temperature parameter.
We further explore the overall performance under different values of $\tau$ with $q=0.5$, as depicted in Figure 3(a). Notably, we observe that a smaller temperature tends to enhance the training process more effectively than a larger one. It is worth noting that excessively high temperatures may exhibit a slightly negative effect on the overall performance.
Dimension of projection head.
We vary the dimension of the projection head $d$ and present the results in Figure 3(b). Our observations indicate that the MFCPL method exhibits robustness across different dimension sizes $d$. However, a noteworthy trend emerges: as we increase the dimension to larger sizes, there is a slight decrease in performance. This finding suggests that a smaller projection head dimension can minimize redundant information and enhance performance.
(a) Temperature $\tau$
(b) Dimension $d$
Figure 4: Performance analysis of MFCPL on three datasets under different values of temperature $\tau$ and projection head dimension $d$ with $q=0.5$.
@@ 5 Conclusion
In this paper, we propose MFCPL, a novel multimodal federated learning framework, breaking through the twin barriers of data heterogeneity and severely missing modalities. Our approach leverages a unique complete prototype methodology, incorporating three key components. We initiate the alignment of complete prototypes at two levels: modality-specific and modality-shared representations, thereby extracting rich semantic representation from multiple modalities. Additionally, we integrate cross-modal alignment to alleviate noise stemming from missing modalities, resulting in significant performance gains. Experiments show that our MFCPL consistently outperforms other state-of-the-art methods, achieving up to $7.69\%$, $5.48\%$ and $8.99\%$ improvements in performance for UCI-HAR, Hateful Memes, and MELD datasets, respectively when $80\%$ of modalities are missing. Thus, MFCPL paves the new way for robust multimodal learning in real-world settings with diverse and incomplete data.

### v2

@@ 4.1.2 Model Architecture
In the case of modalities such as image, text, video and audio, we follow (Feng et al., 2023) to extract features for local training. This method involves deriving features instead of training from scratch using raw data, a choice made to alleviate the computational burden. Specifically, for the visual modality such as image and video, MobileNetV2 (Howard et al., 2017) is employed to extract representations, while textual features are derived using MobileBERT (Sun et al., 2020). For audio data, Mel-frequency cepstral coefficients (MFCCs) (Chen and Rudnicky, 2023) are utilized to extract relevant features. For other modalities, raw data is utilized for training process. Our model architecture is composed of three key modules: the encoder, the fusion layer, and the classifier. For audio, accelerometer, and gyroscope data, the encoder adopts a 1D Conv+RNN architecture. The 1D Conv includes three CNN layers followed by two MLP layers, while the RNN component incorporates a single GRU layer. In contrast, the encoder for images is built with two MLP layers, and for text and video, we employ an RNN architecture with a single GRU layer. The fusion layer is implemented as a self-attention fusion layer, and the classifier is subsequently followed by two MLP layers. The hidden size for GRU layer is set as $128$ for all datasets. Regarding the projection head, we utilize a single MLP layer with the default dimension set to $64$. We provide the detailed model architecture for each dataset in Table 91011.
4.1.3 Implementation Details
@@
Following the methodology outlined in (Feng et al., 2023), we set a uniform missing rate of $q=\{0.5,0.7,0.8,1.0\}$ for each modality based on Bernoulli distribution, creating a scenario that simulates the presence of missing modalities. The Bernoulli distribution with parameter $1-q$ determines the presence, and $q$ determines the absence of each modality for each client. For example, if $q=0.7$, each modality has a $70\%$ probability of being missing and $30\%$ probability of being retained. For each modality $m$, if a sample from the Bernoulli distribution is 00, the modality is retained; if it is $1$, the modality is considered missing. In the case where $q=1.0$, each client will possess only data from one modality while other modalities are missing (cross-modal setting). This approach ensures a uniform missing rate while providing the randomized missing patterns to mimic real-world scenarios. Table 1 illustrates the client distribution across three datasets with different missing rates. For instance, in MELD dataset ($M=3$), there are possible $7$ types of clients: audio, text, video, audio-text, audio-video, text-video, and audio-text-video. Furthermore, to simulate the scenario where clients can access a specified proportion of the missing modality’s data, we randomly sample the amount of data to be missing with the zero-filling data rate $u$. Note that, the missing modality will be filled by zeros as the input. For instance, in the case with a modality missing rate $q=0.5$, clients can access the missing modality’s data with zero-filling data rate $u$ (e.g, $u=0.2$ indicates $20\%$ data in missing modality is absent). The communication round is set to $200$, and the local training epoch is $1$ for all datasets. The default value for the temperature $\tau$ is $0.1$. The batch size is set to $16$, and we employ the SGD (Robbins and Monro, 1951) optimizer with a weight decay of $51e-$ 5, maintaining a consistent learning rate of $0.05$ across all datasets. We set $\alpha_{reg}=1$, $\alpha_{con}=2$, and $\alpha_{align}=0.1$. For the Hateful Memes dataset, we use the AUC score as the evaluation metric. The UAR score serves as the metric for the MELD dataset, and the F1 score is employed as the metric for the UCI-HAR dataset. We repeat the experiment $5$ times and report mean values.
4.1.4 Baselines
@@
For evaluation, we compare our MFCPL with several state-of-the-art FL methods in MFL setting: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FedOpt (Reddi et al., 2020), (4) MOON (Li et al., 2021), prototype-based FL methods: (5) FedProto (Tan et al., 2022b) (with parameter averaging), (6) FedProc (Mu et al., 2023), (7) FPL (Huang et al., 2023), and multimodal federated learning methods: (8) FedIoT (Zhao et al., 2022), (9) FedMSplit (Chen and Zhang, 2022).
(a) UCI-HAR
(b) Hateful Memes
(c) MELD
@@
Figure 5: Illustrative example of performance versus communication rounds three datasets with $q=0.5$.
@@
$q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$ $q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$ $q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$
70.90 69.46 67.50 66.85 54.56 50.09 49.56 47.66 51.22 50.62 50.36 47.46
75.09 74.31 73.63 72.44 53.96 54.92 54.15 52.39 52.65 51.98 50.91 49.53
71.05 70.55 71.68 70.95 51.42 51.58 50.23 49.98 51.94 50.82 50.70 48.88
73.56 70.96 70.41 70.02 51.79 51.17 50.63 50.16 51.98 51.15 50.91 49.82
77.44 75.61 75.19 73.93 56.30 55.69 55.04 53.88 54.71 53.62 52.18 50.26
4.2 Performance Comparison
Table 2 illustrates the performance comparison of our proposed MFCPL with other state-of-the-art methods on three multimodal federated datasets. Our method, MFCPL, consistently outperforms other baselines across various missing rates, highlighting its effectiveness in achieving better generalized global models and supporting clients with missing data. As the missing rate parameter $q$ increases, MFCPL maintains its performance, while other baselines exhibit a significant decrease in accuracy. Taking the UCI-HAR dataset with $q=0.5$ as an illustrative example, our method surpasses the best-performing baseline FPL by a notable margin, showcasing an improvement of $2.28\%$. Similar to the UCI-HAR dataset, MFCPL not only outperforms other methods on different datasets but also maintains its performance as the modality missing rate increases. Notably, when $80\%$ data of modality is missing, our method achieves improvement up to $5.48\%$ and $5.24\%$ in performance on Hateful Memes and MELD datasets, respectively. Furthermore, under a $100\%$ missing rate, where clients possess data from only a single modality, our method demonstrates improvements up to $18.1\%$, $6,73\%$ and $13.46\%$ in performance on UCI-HAR, Hateful Memes and MELD datasets, respectively. These results underscore the robustness and effectiveness of MFCPL by leveraging the complete prototypes to guide the local training in handling missing data scenarios across different client types and datasets. Moreover, MFCPL depicts the key advantages of maintaining the high accuracy even the modality missing rate increase, demonstrating the strong adaptability to scenarios with severe modality absences. Figure 5 depicts the performance curves (performance versus communication rounds) of our MFCPL and SOTA baselines on three datasets, from which we can observe that MFCPL demonstrates a more stable convergence compared to other methods, particularly under non-IID and missing modality conditions.
Furthermore, we explore an alternative practical setting for MFL under missing modalities, wherein clients have access to a portion of the data in the missing modality. Table 3 presents a performance comparison of our proposed method with other state-of-the-art approaches across various data rates $u=\{0.2,0.5,0.7\}$. Similar to the scenario with $100\%$ missing data rate, MFCPL consistently outperforms other baselines across different zero-filling data rates, highlighting its effectiveness in diverse practical settings. Specifically, our proposed scheme achieves the improvements ranging from $1.5\%$ to $4\%$ compared to other baselines across most scenarios. Overall, as the zero-filling data rate increases, all methods across all datasets exhibit a decrease in performance due to the prevalence of zero-filled data. However, MFCPL maintains comparable performance even as the data rate rises, demonstrating the importance of integrating components at both modality-shared and modality-specific levels.
Table 5: Performance analysis of MFCPL on utilizing complete prototypes and unimodal prototypes.
Prototype UCI-HAR
$q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$
Unimodal Prototypes 77.09 74.67 73.34 71.65
Complete Prototypes 77.44 75.61 75.19 73.93
Prototype Hateful Memes
$q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$
Unimodal Prototypes 54.95 54.16 53.68 52.31
Complete Prototypes 56.30 55.69 55.04 53.88
Prototype MELD
$q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$
Unimodal Prototypes 53.32 52.43 50.41 49.14
Complete Prototypes 54.71 53.62 52.18 50.26
@@ 4.3 Comparison to Unimodal Prototypes
4.3.1 Details on MFCPL with Unimodal Prototypes
We utilize the projection head $g_{1}$ to map the unimodal representations $z_{m}$ from encoders $f_{m}$ into shared representation space $\mathbb{R}^{d}$:
$c_{m}=g_{1}(z_{m})\in\mathbb{R}^{d},\forall m\in\{1,2,...,M\}.$ (11)
We define the $k^{th}$ class prototypes for each modality as:
$u^{k}_{m_{i}}=\frac{1}{|S^{k}_{i}|}\sum_{i\in S^{k}_{i}}c_{m_{i}},\forall m\in\{1,2,...,M\},$ (12)
where $S^{k}_{i}$ denotes the subset of $D_{i}$ belonging to $k^{th}$ The unimodal prototypes on the server for each modality are formulated as follows:
$\begin{split}U_{m}^{k}&=\frac{1}{N}\sum_{i\in N}u^{k}_{m_{i}}\in\mathbb{R}^{d},\forall m\in\{1,2,...,M\}\\ U_{m}&=[U_{m}^{1},U_{m}^{2},\dots,U_{m}^{K}],\end{split}$ (13)
where $U_{m}^{k}$ represents the unimodal prototype of each modality $m$ belonging to class $k\in K$. The calibration is conducted between the client type and its corresponding modality’s prototype following the MFCPL framework.
(a) $q=0.5$
(b) $q=0.7$
(c) $q=0.8$
(d) $q=1.0$
Figure 6: Performance analysis of MFCPL on three datasets under different values of temperature $\tau$ with various missing rates.
(a) $q=0.5$
(b) $q=0.7$
(c) $q=0.8$
(d) $q=1.0$
Figure 7: Performance analysis of MFCPL on three datasets under different projection head dimensions $d$ with various missing rates.
4.3.2 Performance Comparison
To assess the efficacy of our proposed complete prototypes, we conducted experiments comparing them with unimodal prototypes on various missing rates, as summarized in Table 5. The results demonstrate the superior performance of utilizing our proposed complete prototypes compared to the unimodal prototypes. Notably, our method exhibits a performance improvement of $1.85\%$, $1.36\%$ and $1.77\%$ on UCI-HAR, Hateful Memes and MELD dataset over the unimodal prototypes with $80\%$ data of modality is missing, respectively. This finding highlights the effectiveness of capturing diverse modality knowledge from complete prototypes, providing valuable guidance for the training process.
4.4 Computation and Communication Scalability
To evaluate the computational cost of our method, we provide a detailed comparison of the computational costs of each additional component (i.e., CMPR, CMPC, and CMA) with the baseline computation cost of FedAvg. Computational complexity is measured in floating-point operations per second (FLOPs) using the FlopCountAnalysis() function from the fvcore library of Meta [^3] https://github.com/facebookresearch/fvcore. Table 6 illustrates the computational cost of each component on the MELD dataset, which includes three modalities (audio, text, and video). As shown in Table 6, the baseline FedAvg incurs a computational cost of 1,421M FLOPs, whereas the combined cost of CMPR, CMPC, and CMA components for prototype computation and alignment is negligible by comparison—approximately 453 times smaller than the CE component. This demonstrates that our proposed design introduces minimal additional computational cost relative to the baseline algorithm (e.g., FedAvg).
Regarding scalability, the computation cost at each client remains constant regardless of the number of federated clients. On the server side, an increase in the number of clients results in more prototypes being transmitted, leading to slightly higher communication costs. However, the size of prototypes is significantly smaller than the model size, making the added communication overhead negligible. Table 7 details the communication costs for each component on the MELD dataset. The results demonstrate that the computational and communication costs of prototype computation are minimal for both the server and clients. Based on these observations, our method demonstrates scalability even when applied to large-scale systems with real-world datasets or a large number of clients. Importantly, the computation cost at each client scales with the size of the local data rather than the total dataset size, ensuring that resource-constrained edge devices can efficiently handle the computation.
Table 6: The computation cost of different components on MELD dataset.
Component FedAvg CMPR CMPC CMA
# FLOPs 1,421M 2.36M 0.39M 0.39M
Table 7: The communication cost (MB) of different components on the MELD dataset. In this experiment, we evaluate with $8$ participating clients in $200$ communication rounds.
Component Model Parameters Local Prototypes Global Prototypes
Cost (MB) $860$ $0.36$ $0.045$
@@ 4.5.1 Effects of Key Components
Table 8: Performance analysis of MFCPL on cross-modal alignment with $q=0.5$.
Case UCI-HAR Hateful Memes MELD
KL 74.86 54.46 52.63
L1 76.62 55.42 53.42
Smooth L1 74.57 55.18 54.95
L2 77.44 56.30 54.71
Table 9: Model architecture used in our experiments for UCI-HAR dataset.
Components Layer Detail
Accelerometer Encoder (1D Conv+RNN) Conv1d(3, 32, kernel_size=5, stride=1, padding=2)
Conv1d(32, 64, kernel_size=5, stride=1, padding=2)
Conv1d(64, 128, kernel_size=5, stride=1, padding=2)
ReLU()
MaxPool1d(kernel_size=2, stride=2)
Dropout(p=0.1)
GRU(128, 128, batch_first=True, dropout=0.1)
Gyroscope Encoder (1D Conv+RNN) Conv1d(3, 32, kernel_size=5, stride=1, padding=2)
Conv1d(32, 64, kernel_size=5, stride=1, padding=2)
Conv1d(64, 128, kernel_size=5, stride=1, padding=2)
ReLU()
MaxPool1d(kernel_size=2, stride=2)
Dropout(p=0.1)
GRU(128, 128, batch_first=True, dropout=0.1)
Self-Attention Linear(128, 512)
Tanh()
Linear(512, 6)
Classifier Head Linear(768, 64)
ReLU()
Dropout(p=0.1)
Linear(64, 6)
Projection Layers Linear(768, 64) for projector
Linear(128, 64) for projector2
Table 10: Model architecture used in our experiments for Hateful Memes dataset.
Components Layer Detail
Image Encoder Linear(1280, 128)
ReLU()
Dropout(p=0.1)
Linear(128, 128)
Text Encoder GRU(512, 128, batch_first=True, dropout=0.1)
Self Attention Fusion Linear(128, 512)
Tanh()
Linear(512, 6)
Classifier Head Linear(768, 64)
ReLU()
Dropout(p=0.1)
Linear(64, 2)
Projection Layers Linear(768, 64) for projector
Linear(128, 64) for projector2
Table 11: Model architecture used in our experiments for MELD dataset.
Components Layer Detail
Audio Encoder (1D Conv+RNN) Conv1d(80, 32, kernel_size=5, stride=1, padding=2)
Conv1d(32, 64, kernel_size=5, stride=1, padding=2)
Conv1d(64, 128, kernel_size=5, stride=1, padding=2)
ReLU()
MaxPool1d(kernel_size=2, stride=2)
Dropout(p=0.3)
GRU(128, 128, batch_first=True, dropout=0.3)
Text Encoder GRU(512, 128, batch_first=True, dropout=0.3)
Video Encoder GRU(1280, 128, batch_first=True, dropout=0.3)
Self-Attention Fusion Linear(128, 512)
Tanh()
Linear(512, 6)
Classifier Head Linear(768, 64)
ReLU()
Dropout(p=0.3)
Linear(64, 4)
Projection Layers Linear(768, 64) for projector
Linear(128, 64) for projector2
4.5.2 Temperature parameter
We further explore the overall performance under different values of $\tau$ with different missing rates, as depicted in Figure 6. Notably, we observe that a smaller temperature tends to enhance the training process more effectively than a larger one. It is worth noting that excessively high temperatures may exhibit a slightly negative effect on the overall performance across various missing rates.
4.5.3 Dimension of projection head
We vary the dimension of the projection head $d$ and present the results in Figure 7. Our observations indicate that the MFCPL method exhibits robustness across different dimension sizes $d$. However, a noteworthy trend emerges: as we increase the dimension to larger sizes, there is a slight decrease in performance. This finding suggests that a smaller projection head dimension can minimize redundant information and enhance performance. It is intriguing to note that even a small projected space dimension ($d=16$) can yield better performance than a larger dimension ($d=128$). This highlights the effectiveness of even a compact projection space in preserving correlations between different representations from various modalities. For instance, in the Hateful Memes dataset, where the original feature space for image and text is $[1280,512]$, such small projection spaces prove to be beneficial.
4.5.4 Ablation study on cross-modal alignment loss
In this experiment, we delve into the examination of cross-modal alignment loss using various types of loss functions across three datasets, as depicted in Table 8, including Kullback-Leibler (KL) Divergence (Kullback, 1997), L1, and smooth L1 loss functions. Our results illustrate that the utilizing L2 loss (MSE loss) predominantly yields superior performance across most datasets. However, it is noteworthy that for datasets like MELD or UCI-HAR, L1 and smooth L1 loss functions also achieve notable performance, such as with smooth L1 loss on the MELD dataset. This observation underscores the adaptability of our cross-modal alignment approach to different loss functions, showcasing its efficacy across varied functions and mitigate the difference of multiple modalities.
@@ 5 Conclusion
In this paper, we propose MFCPL, a novel multimodal federated learning framework, breaking through the twin barriers of data heterogeneity and severely missing modalities. Our approach leverages a unique complete prototype methodology, incorporating three key components. We initiate the alignment of complete prototypes at two levels: modality-specific and modality-shared representations, thereby extracting rich semantic representation from multiple modalities. Additionally, we integrate cross-modal alignment to alleviate noise stemming from missing modalities, resulting in significant performance gains. Experiments show that our MFCPL consistently outperforms other state-of-the-art methods under severely missing modality scenarios, achieving up to $7.69\%$, $5.48\%$ and $5.24\%$ improvements in performance for UCI-HAR, Hateful Memes, and MELD datasets, respectively, when $80\%$ data of modality is missing. Thus, MFCPL paves the new way for robust multimodal learning in real-world settings with diverse and incomplete data.